OpenStudy (melodycrazysouthern):

(2.8 × 109) ÷ (4 × 103)

OpenStudy (anonymous):

use the order of operations to solve this :)

OpenStudy (melodycrazysouthern):

ow cool what is order or operation :(

OpenStudy (anonymous):

PEMDAS P: Parenthesis E: Exponent M: Multiplication D: Division A: Addition S: Subtraction

Parth (parthkohli):

Seriously?

OpenStudy (melodycrazysouthern):

oww ok

OpenStudy (anonymous):

So... u guys aint gonna give the answer to them... i gotchu melody...

OpenStudy (anonymous):

You can use: 'Please Excuse My Dear Aunt Sally' to remember :)

Parth (parthkohli):

^8\[\frac{2.8 \times 10^9}{4\times10^3} = \frac{28 \times 10^8}{4 \times 10^3} = \frac{28}{4}\times \frac{10^8}{10^3} = 7 \times 10^5\]

Parth (parthkohli):

So you can see, there's really nothing here about the order of operations.

OpenStudy (steve816):

@RainbowBrony555 Please End My Day At School

OpenStudy (melodycrazysouthern):

lol haha

OpenStudy (steve816):

@BweadedChicken LOL You're going to get suspended if you don't delete that xD

OpenStudy (anonymous):

Bro its ok xD PeerAnswer is better, i just come here like once in a blue moon anyways

OpenStudy (anonymous):

Well anyway, got the answer? :)

OpenStudy (anonymous):

Well damn, thanks for the medals guys, i feel special

OpenStudy (melodycrazysouthern):

lol dont get expended BweadedChicken

OpenStudy (anonymous):

expended? xD come join the mathematics chat

OpenStudy (melodycrazysouthern):

lol

OpenStudy (melodycrazysouthern):

so boring

OpenStudy (melodycrazysouthern):

boring

OpenStudy (anonymous):

The answer is... OVER 9000!!!!!!